INTERNAL ENERGY DISTRIBUTION OF CH_{3}O (X^{2}E) FROM PHOTODISSOCIATION OF METHYL NI-TRATE

\underline{ERIKA L. DERRO}, CRAIG MURRAY, MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104; and MARK D. MARSHALL, Department of Chemistry, Amherst College, Amherst, MA 01002.

The methoxy (CH_{3}O) radical has attracted a great deal of attention in the literature due to its unique spectroscopy, as well as its importance as a reactive intermediate in combustion and atmospheric chemistry. In this study, we employ CH_{3}O as a dynamical probe of the photodissociation dynamics of methyl nitrate, CH_{3}ONO_{2}, at 193 nm. Rotationally-resolved spectra of the CH_{3}O photoproduct were recorded in the A^{2}A_{1} − X^{2}E system under both nascent and jet-cooled conditions using laser-induced fluorescence (LIF) spectroscopy. CH_{3}O A^{2}A_{1} − X^{2}E transitions could be simulated effectively using known spectroscopic constants at the lowest temperatures achieved under jet-cooled conditions. Spectra recorded at warmer temperatures are being used to refine reported spectroscopic constants. Under nascent conditions, transitions were observed from the vibrationless level of CH_{3}O (X^{2}E) as well as with C−O stretch vibrational excitation, and the vibrational branching ratio is determined. The rotational excitation of these vibrational bands was fit to a Boltzmann distribution to extract nascent rotational temperatures. The experimental data indicate that the CH_{3}O fragment is produced with minimal internal energy (1% of the available energy) following photodissociation of CH_{3}ONO_{2} at 193 nm.